Circular No. 7396

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/2000 G1 (LINEAR) F. Shelly, for the Lincoln Near-Earth Asteroid Research project, re-

ports, in connection with the discovery yesterday of a fast-moving object, that Lisa Brown-Manguso noticed that the object showed clear cometary activity. The following CCD astrometry is available:

2000 UT α_{2000} δ_{2000} m_{1} Observer
Apr. 7.45368 18^{h}06^{m}50.^{s}72 −7^{o}26′35′′.3 17.9 Brown-Manguso
7.48656 18 06 50.91 −7 25 13.7 18.3  _{”}
8.09218 18 07 02.04 −7 00 18.9 17.6 Tichá
8.09337 18 07 02.08 −7 00 15.3  _{”}
8.09478 18 07 02.04 −7 00 11.9  _{”}
8.09581 18 07 01.95 −7 00 07.8  _{”}
8.43066 18 07 07.01 −6 46 31.8 17.5 Hug
8.43301 18 07 07.01 −6 46 26.6 17.2  _{”}
8.43635 18 07 06.97 −6 46 20.1 17.9  _{”}

L. Brown-Manguso, M. Blythe, F. Shelly, M. Bezpalko, M. Elowitz and R. Huber (LINEAR). 1.0-m f/2.15 reflector. Measurers J. Stuart, H. Viggh, R. Sayer and J. B. Evans. Wide diffuse tail in p.a. 270^{o}.

J. Tichá, M. Tichý and Z. Moravec (Kleť). 0.57-m f/5.2 reflector. Measurer

Tichý. Object diffuse, coma diameter 13′′. G. Hug (Eskridge). 0.3-m Schmidt-Cassegrain. Coma ∼ 1′ in diameter

with sharp condensation (but not starlike), short tail in p.a. 275^{o}.

Parabolic orbital elements follow. It is very likely that the comet is of short period (with current Δ ∼ 0.24 AU).

T = 2000 Mar. 9.02 TT  ω = 341.^{o}44
Ω = 194.56 }2000.0
q = 1.0112 AU  i = 13.77
2000TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Apr. 6 18^{h}06.^{m}29 − 8^{o}29.^{′}1 0.306 1.117 104.^{o}4 60.^{o}2 16.9
11 18 07.50 − 5 08.4 0.338 1.155 108.4 55.4 17.3
16 18 07.01 − 2 25.4 0.371 1.197 112.7 50.7 17.6
21 18 05.07 − 0 12.8 0.403 1.242 117.2 46.0 18.0
26 18 01.87 + 1 34.5 0.436 1.290 121.8 41.5 18.3
May 1 17 57.60 + 2 59.5 0.470 1.341 126.5 37.2 18.6
6 17 52.46 + 4 04.5 0.506 1.395 131.2 33.0 19.0
2000 April 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Brian G. Marsden